# Quenched shear stresses and shear moduli in disordered elastic networks

### v1

@@ Abstract
We consider theoretically and numerically disordered elastic networks with annealed or quenched connectivity matrices (CMs) focusing on the quenched shear stresses induced by the spontaneous symmetry breaking between different CMs. By rewriting the stress-fluctuation formalism relation for the shear modulus it is argued that theses quenched stresses give an important contribution $\mu_{1}$ to the finite shear modulus $\mu_{\mathrm{q}}=\mu_{1}+\mu_{\mathrm{a}}$ for quenched CMs. $\mu_{\mathrm{a}}$ is a simple average over all states and CMs which for equilibrium distributions of CMs is equivalent to the shear modulus of annealed networks. Interestingly, $\mu_{\mathrm{a}}$ may be strongly negative for out-of-equilibrium distributions as shown by considering a temperature quench and a scalar active two-temperature model.
Introduction.
@@
It is generally accepted that the universe started with an extremely simple and highly symmetrical quantum state Gell-Mann (1995) which due to spontaneous symmetry-breaking fluctuations of all sorts Gell-Mann (1995); Anderson (1964); Ashcroft and Mermin (1976); Chaikin and Lubensky (1995); foo (a) say, between matter and anti-matter dominated worlds, unfolded to its present state. Broken symmetries may become permanently frozen in time, e.g., our universe should remain matter rather than anti-matter dominated, but they may also be only transiently frozen being again washed away by subsequent fluctuations after some characteristic time $t_{\star}$. Broken (permanent or transient) symmetries do not only play a role for crystalline solids Anderson (1964); Ashcroft and Mermin (1976) but are also crucial for complex matter systems without “broken continuous symmetries” but frozen local heterogeneities Götze (2009). We investigate in the present work amorphous solids formed by ideal spring networks with randomly frozen “connectivity matrices” (CMs) describing the interactions between the connected particles (“vertices”) foo (b). Due to the quenched disorder we need to be more precise concerning the averaging procedure for fluctuations George et al. (2021) being important for the determination of second derivatives of the free energy Chaikin and Lubensky (1995), especially for the shear stress relaxation modulus $\mu$ computed by means of the stress-fluctuation formalism Squire et al. (1969); Wittmer et al. (2013, 2016); George et al. (2021). We thus analyze various “moments” of the instantaneous shear stress $\hat{\tau}(c,k)$ with $k$ denoting the thermodynamic state of a CM $c$. We denote, e.g., by $\tau_{\mathrm{q}}(c)\equiv\left<\hat{\tau}(c,k)\right>_{k}$ the $k$-averaged quenched shear stress for a given $c$. The (rescaled) variance $\mu_{1}$ of these quenched shear stresses is shown to yield an important contribution to the shear modulus $\mu$.
We begin by stating the central theoretical relations assuming permanent CMs ($t_{\star}=\infty$) distributed according to an equilibrium procedure. As illustrated using a simple computational model network with annealed and quenched disorder we then generalize these relations into the time domain by allowing the reorganization of the CMs, i.e. $t_{\star}(\omega)$ becomes finite with $\omega$ being a convenient tuning parameter. This is further generalized by considering a temperature quench with an initial out-of-equilibrium distribution of CMs being slowly allowed to relax and “scalar active” networks Grosberg and Joanny (2015); Elismaili et al. (2020) with different temperatures for the vertices and the annealed CMs.
Theoretical considerations.
@@
Let $\beta=1/T$ denote the inverse temperature (in units of Boltzmann’s constant), $V$ the system volume and $E(c,k)$ the energy of a state $k$ for a CM $c$. Following Refs. Wittmer et al. (2013, 2016); George et al. (2021) we expand $E(c,k)$ assuming a canonical transformation George et al. (2021) of the positions and momenta of the network vertices under an infinitessimal affine simple shear strain $\gamma$. The instantaneous shear stress $\hat{\tau}(c,k)$ and the instantaneous affine shear modulus $\hat{\mu}_{\mathrm{A}}(c,k)$ are defined as first and second functional derivatives of $E(c,k)/V$ with respect to this transformation Wittmer et al. (2016). While the frozen shear stresses $\tau_{\mathrm{q}}(c)\equiv\left<\hat{\tau}(c,k)\right>_{k}$ are in general finite, the ensemble of all networks has by symmetry an average shear stress
@@
being all $\geq 0$. Due to Eq. (1) the (rescaled) variance $\mu_{1}$ is a measure of the symmetry breaking associated with the quenched shear stresses. $\mu_{\mathrm{F}}$ characterizes the typical shear-stress fluctuations for each $c$ and $\mu_{\mathrm{A}}$ the $k$- and $c$-averaged “affine shear modulus”. According to the stress-fluctuation formalism Squire et al. (1969); Wittmer et al. (2013, 2016); George et al. (2021) $\mu(c)\geq 0$ is the (thermodynamic) shear modulus of each CM $c$. Let us rewrite the $c$-averaged shear modulus $\mu$ as
@@
What is the meaning of $\mu_{\mathrm{a}}$? We denote by $l$ all combined states lumping together all $(c,k)$. Reminding also Eq. (1), the “simple average” Allen and Tildesley (2017); George et al. (2021)
@@
corresponds to the formal application of the stress-fluctuation formalism disregarding the assumed symmetry breaking. To be a thermodynamically meaningful modulus not only the states $k$ of the given CMs $c$ must be at thermal equilibrium, however, but also the CMs. To see why $\mu_{\mathrm{a}}\geq 0$ for equilibrium CM-distributions let us assume that the broken symmetry becomes lifted by reversible reorganizations of the CMs for $t\gg t_{\star}$. Using Eq. (1) implies that $\mu_{\mathrm{a}}=\mu-\mu_{1}=\mu\geq 0$ is the shear modulus of annealed networks. Being a simple average $\mu_{\mathrm{a}}$ does not change if evaluated for permanently quenched CMs of same distribution albeit it differs then from the true thermodynamic shear modulus $\mu_{\mathrm{q}}=\mu_{\mathrm{a}}+\mu_{1}$ of strictly quenched CMs with finite $\mu_{1}$. ($\mu_{\mathrm{q}}$ is often used instead of $\mu$ for strictly quenched CMs.) Interestingly,
$\mu_{\mathrm{a}}\geq 0\Leftrightarrow\mu_{\mathrm{A}}\geq\mu_{2}\Leftrightarrow\mu_{\mathrm{q}}\geq\mu_{1}.$ (10)
These inequalities become equalities if the annealed system behaves for $t\gg t_{\star}$ like a liquid, i.e. $\mu_{1}=\mu=0$. For arbitrarily slowly quenched colloidal or polymer glasses kept at thermal equilibrium one thus expects Wittmer et al. (2016); George et al. (2021)
$\mu_{\mathrm{a}}=\mu_{\mathrm{A}}-\mu_{2}=0\mbox{ and }\mu_{\mathrm{q}}=\mu_{1}.$ (11)
Figure 1: Annealed elastic network with $0.1\leq R_{s}\leq 2$ and $R_{s}\right>_{s}=1$ at $T=0.1$: (a) Snapshot of a small subvolume comprising about 500 vertices. Red lines indicate compressed springs ($r<R_{s}$), blue lines extended springs. (b) Normalized distribution $p(R_{s})$.
Numerical model.
@@
We use for illustration purposes an extremely simple two-dimensional model elastic network where $n$ vertices of mass $m$ are connected by ideal springs $s$ of energy $\frac{K}{2}(r-R_{s})^{2}$ with identical spring constants $K$ but a specific reference length $R_{s}$ for each spring. We assume $m=1$, $K=10$ and $R_{s}\right>_{s}=1$ for the first moment of the $R_{s}$-distribution. The network topology is set up using a regular hexagonal lattice of springs where each vertex is connected by identical springs of $R_{s}=1$ to its six next neighbors. At variance to the transient spring networks considered in Ref. Wittmer et al. (2016) the springs remain permanently connected to the same vertices, i.e. with this respect all CMs are quenched and identical. We use below $n=10^{4}$ vertices contained in periodic rectangular simulation boxes of linear lengths $L_{x}=100$ and $L_{y}=L_{x}\sqrt{3}/2$. The networks are sampled by means of molecular dynamics (MD) using a Langevin thermostat with a large friction constant $\zeta=1$. The standard temperature is $T=0.1$. For crystalline networks without disorder $\tau_{\mathrm{q}}(c)=0$ by symmetry for all $c$ and $T$ and, hence, $\mu_{1}=0$ and $\mu=\mu_{\mathrm{a}}$.
@@
Annealed and quenched disorder enters by allowing $R_{s}$ to fluctuate under the additional constraint that $0.1\leq R_{s}\leq 2$. This is done by means of a Monte Carlo (MC) procedure where the reference lengths of two (randomly chosen) neighboring springs are slightly changed subject to a standard Metropolis criterion Allen and Tildesley (2017) and that both new lengths are within the indicated bounds and their sum remains constant. (This may be seen as a simple model for branched wormlike micelles with imposed number of branching points and total branch length Cates and Candau (1990).) The algorithm thus mixes MD moves for the vertices at a given CM with MC steps altering the Hamiltonian and thus the CM foo (b). Time is measured by the MD procedure where we set $\delta t=10^{-3}$ for the Verlet time increment Allen and Tildesley (2017). After each $\delta t$ we attempt for a small fraction of pairs of springs an MC hopping move to change their reference lengths. The reorganization frequency $\omega$ indicated below is proportional to this small fraction: $\omega=1$ means that a spring is attempted on average once per time unit. Results obtained for $\omega=0.01$ are shown in Fig. 1. As can be seen in panel (a), the network becomes very disordered and appears macroscopically homogeneous and isotropic. This visual impression can be corroborated using standard pair correlation functions Allen and Tildesley (2017) (not shown). Panel (b) shows the corresponding probability distribution for $R_{s}$. We thus obtain a large ensemble of $N_{\mathrm{c}}=500$ independent configurations $c$ with an equilibrium CM-distribution. Assuming ergodicity for each CM we obtain $k$-averages by means of time-averages Allen and Tildesley (2017). For instance, we estimate the $k$-average $\tau_{\mathrm{q}}(c)$ by computing the time-average $\tau_{\mathrm{q}}(c,t)$ of the instantaneous shear stresses $\hat{\tau}$ over a “sampling time” $t$. The above-introduced moments thus become apriori $t$-dependent as emphasized by an additional argument $t$ foo (c). We remind that for stationary systems the dynamical shear modulus $\mu(t)$ can be shown to be a smoothing function of the shear stress relaxation modulus $G(t)$ George et al. (2021); foo (d). Data are sampled either for strictly permanent networks ($\omega=0$) or by very gradually reorganizing the CMs at small finite $\omega$. The behavior of permanent CMs is then obtained from the large-$t$ and small-$\omega$ asymptotic limits.

### v2

# On inequalities of shear modulus contributions in disordered elastic bodies

@@ Abstract
We investigate generic inequalities of various contributions to the shear modulus $\mu$ in ensembles of amorphous elastic bodies. We focus first on a simple elastic network model with connectivity matrices (CMs) which are either annealed or quenched, at or out of equilibrium. The stress-fluctuation formalism relation for $\mu$ is rewritten as $\mu=\mu_{1}+\mu_{\mathrm{a}}$ with $\mu_{1}\geq 0$ characterizing the variance of the quenched shear stresses and $\mu_{\mathrm{a}}$ being a simple average over all states and CMs. For equilibrium CM-distributions $\mu_{\mathrm{a}}$ becomes equivalent to the shear modulus of annealed systems, i.e. $\mu_{\mathrm{a}}\geq 0$, while more generally $\mu_{\mathrm{a}}$ may become strongly negative as shown by considering a temperature quench and a scalar active two-temperature model. Consistent relations are also found for glass-forming colloids where $\mu-\mu_{1}=\mu_{\mathrm{a}}=0$ for equilibrium ensembles, i.e. $\mu$ is set by the quenched shear stresses, while $\mu_{\mathrm{a}}$ becomes again negative otherwise.
I Introduction
Figure 1: Elastic ideal spring network with annealed reference lengths $R_{s}$ of the springs imposing $0.1\leq R_{s}\leq 2$ and $R_{s}\right>_{s}=1$ for a temperature $T=0.1$: (a) Snapshot of a small subvolume comprising about 500 vertices. Red lines indicate compressed springs ($r<R_{s}$), blue lines extended springs. (b) Normalized distribution $p(R_{s})$.
General context and goal.
@@ I Introduction
Broken symmetries Anderson (1964); Chaikin and Lubensky (1995); foo (a) may become permanently frozen in time (“$t_{\star}=\infty$”) but they may also be only transient being annealed after some finite characteristic relaxation time $t_{\star}$. Permanent or transiently broken symmetries do not only play a role for crystalline solids Anderson (1964); Ashcroft and Mermin (1976) but are also crucial for complex matter systems without “broken continuous symmetries” but frozen local heterogeneities Dhont (2001); Barrat (2006); van Workum and de Pablo (2003); Yoshimoto et al. (2004); Tsamados et al. (2009); Mizuno et al. (2013); Puosi et al. (2014); Bouzid and Gado (2018); Zhang et al. (2019); Wittmer et al. (2023); Semenov and Baschnagel (2024). As sketched in Fig. 1, we investigate in the present study amorphous solids formed by ideal spring networks with randomly frozen “connectivity matrices” (CMs) describing the interactions between the connected particles (“vertices”) foo (b) and, more briefly, standard colloidal glasses created by a temperature quench from equilibrated high-temperature configurations Dhont (2001); Wittmer et al. (2013); Ninarello et al. (2017). Due to the quenched disorder one naturally needs to be more precise concerning the averaging procedure of fluctuations George et al. (2021) being crucial for the determination of second derivatives of the free energy Chaikin and Lubensky (1995). Specifically, we shall focus on the shear modulus $\mu$ of amorphous solids computed by means of the stress-fluctuation formalism Rowlinson (1959); Squire et al. (1969); Lutsko (1989); Barrat et al. (1988); Wittmer et al. (2002); Maloney and Lemaître (2004); Yoshimoto et al. (2005); Schnell et al. (2011); Wittmer et al. (2013, 2016); Procaccia et al. (2016); Klochko et al. (2021); George et al. (2021). We thus analyze various “moments” of the instantaneous shear stress $\hat{\tau}(c,k)$ with $k$ denoting the thermodynamic state of a CM $c$. For instance, we denote by $\tau_{\mathrm{q}}(c)\equiv\langle\hat{\tau}(c,k)\rangle_{k}$ the $k$-averaged quenched shear stress for a given $c$. Consistently with the more specific cases discussed elsewhere Cates et al. (1998); O’Hern et al. (2002, 2003); Zaccone and Terentjev (2013); Wittmer et al. (2016); Cui et al. (2019); Bouzid and Gado (2018); Zhang et al. (2019); Klochko et al. (2021), the rescaled variance $\mu_{1}$ of these quenched shear stresses is quite generally shown to yield an important contribution to the shear modulus $\mu=\mu_{1}+\ldots$ This is especially relevant for quenched liquid-type glass-formers under equilibrium conditions (as properly defined in Sec. VI) for which we shall demonstrate that
$0=\mu-\mu_{1}\mbox{ with }\mu_{1}=\beta V\left[\langle\tau_{\mathrm{q}}(c)^{2}\rangle_{c}-\langle\tau_{\mathrm{q}}(c)\rangle_{c}^{2}\right]$ (1)
corroborating a recently made suggestion Wittmer et al. (2016). ($\beta=1/T$ stands for the inverse of the temperature $T$, $V$ for the $d$-dimensional volume.)
Outline.
We begin in Sec. II by reminding the stress-fluctuation relations for the shear modulus and state then general inequalities assuming permanent CMs ($t_{\star}=\infty$) taken from an equilibrium distribution. The simple elastic network model shown in Fig. 1 is described in Sec. III. The static relations of Sec. II are generalized into the time domain foo (c) by allowing the reorganization of the CMs of the networks, i.e. $t_{\star}(\omega)$ becomes finite with the imposed network reorganization frequency $\omega$ being a convenient tuning parameter. Computational results obtained for this model are presented in Sec. IV for CMs at equilibrium and in Sec. V for out-of-equilibrium CMs. We consider especially an initially out-of-equilibrium CM-distribution caused by a temperature jump and “scalar active” networks Grosberg and Joanny (2015); Elismaili et al. (2020) with different temperatures for the network vertices and the annealed CMs. In Sec. VI we turn finally to quenched colloidal glasses and confirm that Eq. (1) must in general hold for quenched ensembles kept at equilibrium. We conclude in Sec. VII with an outlook on a similar relation for the microscopic shear modulus being relevant, e.g., for reversible physical gels Wittmer et al. (2016); Bouzid and Gado (2018); Zhang et al. (2019) or polymer melts Rubinstein and Colby (2003).
II Theoretical considerations
Stress-fluctuation formalism.
Having first been used by Rowlinson for the compression modulus of liquids Rowlinson (1959) and later generalized by Squire et al. Squire et al. (1969) for solid bodies, the stress-fluctuation formalism provides a convenient route to calculating the (isothermic) elastic constants in computer simulations Tadmor and Miller (2011). It has thus been widely used in the past both for macroscopic elastic moduli Squire et al. (1969); Lutsko (1989); Barrat et al. (1988); Wittmer et al. (2002); Maloney and Lemaître (2004); Yoshimoto et al. (2005); Schnell et al. (2011); Wittmer et al. (2013, 2016); Procaccia et al. (2016); Klochko et al. (2021); George et al. (2021) as for local mechanical properties Lutsko (1989); Barrat (2006); van Workum and de Pablo (2003); Yoshimoto et al. (2004); Tsamados et al. (2009); Mizuno et al. (2013); Puosi et al. (2014); Wittmer et al. (2023); Semenov and Baschnagel (2024). We note that the relations given in the literature depend somewhat on whether the elastic moduli are defined with respect to an unstressed reference state Squire et al. (1969); Tadmor and Miller (2011) or with respect to an in general prestressed reference Rowlinson (1959). The difference between both definitions is described by so-called “Birch coefficients” as discussed elsewhere Birch (1938); Wallace (1970); Schnell et al. (2011). We take here the latter more physical definition where for gases and liquids the shear modulus strictly vanishes and the compression modulus is consistent with Rowlinson’s formula Rowlinson (1959); Allen and Tildesley (2017); Wittmer et al. (2013).
Shear modulus of amorphous solids.
We summarize here the salient relations for the macroscopic shear modulus $\mu$ relevant for ensembles of quenched amorphous bodies. Let $E(c,k)$ denote the energy for a state $k$ of a CM $c$ at a given volume $V$ and shear strain $\gamma$. Following Refs. Wittmer et al. (2016); Procaccia et al. (2016); George et al. (2021) we expand $E(c,k)$ assuming a canonical transformation George et al. (2021) of the positions and momenta of the particles of the system under an infinitessimal affine simple shear strain $\gamma$ in, say, the $xy$-plane of the system. The instantaneous shear stress $\hat{\tau}(c,k)$ and the instantaneous affine shear modulus $\hat{\mu}_{\mathrm{A}}(c,k)$ are then defined by
$\hat{\tau}(c,k)\equiv\frac{\delta E(c,k)/V}{\delta\gamma}\mbox{ and }\hat{\mu}_{\mathrm{A}}(c,k)\equiv\frac{\delta^{2}E(c,k)/V}{\delta\gamma^{2}}$ (2)
@@ II Theoretical considerations
as the first and the second functional derivatives with respect to this canonical transformation taken at $\gamma=0$ for constant $V$ Wittmer et al. (2016). While the frozen shear stresses $\tau_{\mathrm{q}}(c)\equiv\left<\hat{\tau}(c,k)\right>_{k}$ are in general finite, the $c$-ensemble has by symmetry an average shear stress
@@
being all $\geq 0$. As shown elsewhere Wittmer et al. (2013); Procaccia et al. (2016), the indicated shear modulus $\mu(c)$ for each $c$ is consistent with the thermodynamic definition
$\mu(c)=\frac{\partial^{2}F(T,V,\gamma,c)/V}{\partial\gamma^{2}},$ (9)
i.e. the second derivative of the free energy $F(T,V,\gamma,c)$ taken at $\gamma=0$ at constant $T$, $V$ and $c$. $\mu_{\mathrm{F}}$ characterizes the typical shear-stress fluctuations for each $c$ and $\mu_{\mathrm{A}}$ the $k$- and $c$-averaged “affine shear modulus”. Due to Eq. (3) the (rescaled) variance $\mu_{1}$ is a measure of the symmetry breaking associated with the quenched shear stresses. This is mathematically manifested by the fact that while the $k$- and $c$-averages commute for the “simple averages” Allen and Tildesley (2017) $\mu_{2}$ and $\mu_{\mathrm{A}}$, this is not the case for the “fluctuation” $\mu_{1}$ since in general
$\langle\langle\hat{\tau}(c,k)\rangle^{2}_{k}\rangle_{c}\neq\langle\langle\hat{\tau}(c,k)\rangle^{2}_{c}\rangle_{k}=\langle\langle\hat{\tau}(c,k)\rangle_{k}\rangle_{c}^{2}=0.$ (10)
Inequalities.
Being a thermodynamic susceptibility $\mu(c)\geq 0$ must hold for each CM $c$ and, hence, also $\mu\geq 0$ for its $c$-average. Let us rewrite the $c$-average as
@@
What is the meaning of $\mu_{\mathrm{a}}$? Let us denote by $l$ all combined states lumping together all $(c,k)$. Reminding also Eq. (3), the “simple average” Allen and Tildesley (2017); George et al. (2021)
@@
would correspond to the stress-fluctuation formula for the shear modulus if the symmetry breaking between different CMs could be ignored, i.e., to be a thermodynamically meaningful modulus not only the states $k$ of the given CMs $c$ must be at thermal equilibrium but also the CMs. To see why $\mu_{\mathrm{a}}\geq 0$ for equilibrium CM-distributions let us assume that the broken symmetry becomes lifted by reversible reorganizations of the CMs for $t\gg t_{\star}$ assuming (temporarily) a finite $t_{\star}$. Using Eq. (3) implies that $\mu_{\mathrm{a}}=\mu-\mu_{1}=\mu$ is the shear modulus of the annealed CMs and, hence, $\mu_{\mathrm{a}}\geq 0$. Interestingly, being a simple average, the expectation value of $\mu_{\mathrm{a}}$ does not change if evaluated for permanently quenched CMs of the same CM-distribution. However, for quenched CMs with $\mu_{1}>0$ it then differs from the shear modulus $\mu_{\mathrm{q}}$ due to the inequality $\mu_{\mathrm{q}}=\mu_{\mathrm{a}}+\mu_{1}>\mu_{\mathrm{a}}$. (Instead of $\mu$ we often use below $\mu_{\mathrm{q}}$ for quenched CMs with $t_{\star}=\infty$.) Interestingly, the above argument implies
$\mu_{\mathrm{a}}\geq 0\Leftrightarrow\mu_{\mathrm{A}}\geq\mu_{2}\Leftrightarrow\mu_{\mathrm{q}}\geq\mu_{1}$ (14)
if the (quenched) CMs are taken from an equilibrium distribution. This is the most central point we want to highlight in this work. The special limit where the above inequalities reduce to Eq. (1) will be further investigated in Sec. VI. Moreover, while Eq. (14) holds for equilibrium CM-distributions it may be violated for more general distributions as shown in Sec. V. We emphasize finally that in the above argument leading to Eq. (14) the specific Hamiltonian was irrelevant and there was no need to specify $\hat{\tau}$ and $\hat{\mu}_{\mathrm{A}}$ defined by Eq. (2). See Refs. Squire et al. (1969); Barrat et al. (1988); Yoshimoto et al. (2005); Wittmer et al. (2013, 2016); Procaccia et al. (2016); George et al. (2021); Cui et al. (2019) for specific cases.
III Simple elastic network model
Springs and topology.
@@ III Simple elastic network model
We use for illustration purposes a simple two-dimensional model elastic network where $n$ vertices of mass $m$ are connected by ideal springs $s$ of energy $\frac{K}{2}(r-R_{s})^{2}$ with identical spring constants $K$ but a specific reference length $R_{s}$ for each spring. We assume $m=1$, $K=10$ and $R_{s}\right>_{s}=1$ for the first moment of the $R_{s}$-distribution for each CM $c$. The network topology is set up using a regular hexagonal lattice of springs where each vertex is connected by identical springs of $R_{s}=1$ to its six next neighbors. At variance to the transient spring networks considered in Ref. Wittmer et al. (2016), the springs remain permanently connected to the same vertices, i.e. in this respect all CMs are quenched and identical. We use below $n=10^{4}$ vertices contained in periodic rectangular simulation boxes. The networks are sampled by means of molecular dynamics (MD) using a Langevin thermostat Allen and Tildesley (2017). For crystalline networks without disorder $\tau_{\mathrm{q}}(c)=0$ by symmetry for all $c$ and $T$ and, hence, $\mu_{1}=0$ and $\mu=\mu_{\mathrm{a}}$.
Annealed and quenched disorder.
@@
Annealed and quenched disorder enters since the reference lengths may fluctuate under the additional constraint that $0.1\leq R_{s}\leq 2$. This is done by means of a Monte Carlo (MC) procedure where the reference lengths of two (randomly chosen) neighboring springs are slightly changed subject to a standard Metropolis criterion Allen and Tildesley (2017) and that both new lengths are within the indicated bounds and their sum remains constant. (This may be seen as a simple model for branched wormlike micelles with imposed number of branching points and total branch length Cates and Candau (1990).) The algorithm thus mixes MD moves for the vertices at a given CM with MC steps altering the Hamiltonian and thus the CM foo (b). Time is measured by the MD procedure where we set $\delta t=10^{-3}$ for the Verlet time increment Allen and Tildesley (2017). After each $\delta t$ we attempt for a small fraction of pairs of springs an MC hopping move to change their reference lengths. The reorganization frequency $\omega$ indicated below is proportional to this small fraction: $\omega=1$ means that a spring is attempted on average once per time unit.
Configurations and data sampling.
Results obtained for $\omega=0.01$ and running the systems over $t=10^{5}$ are shown in Fig. 1. As can be seen in panel (a), the network becomes very disordered and appears macroscopically homogeneous and isotropic. (This visual impression can be corroborated using standard pair correlation functions.) Panel (b) shows the corresponding probability distribution for $R_{s}$. We thus obtain a large ensemble of $N_{\mathrm{c}}=500$ independent configurations $c$ with an equilibrium CM-distribution. Assuming ergodicity for each CM $k$-averages are replaced by time-averages Allen and Tildesley (2017). For instance, we estimate the $k$-average $\tau_{\mathrm{q}}(c)$ by computing the time-average $\tau_{\mathrm{q}}(c,t)$ of the instantaneous shear stresses $\hat{\tau}$ over a “sampling time” $t$. The above-introduced moments thus become apriori $t$-dependent as emphasized by an additional argument $t$ foo (c). We remind that for stationary systems the dynamical shear modulus $\mu(t)$ is a smoothing function George et al. (2021); foo (d) of the shear stress relaxation modulus $G(t)$ Rubinstein and Colby (2003). Data are sampled either for strictly permanent networks ($\omega=0$) or by gradually reorganizing the CMs at small finite $\omega$ as in our work on transient self-assembled spring networks Wittmer et al. (2016); George et al. (2021). The behavior of permanent CMs is then obtained from the small-$\omega$ asymptotic limit.
IV Elastic networks at equilibrium